OpenStudy (anonymous):

FACOTR OF 121X^2-22X+1

OpenStudy (anonymous):

REMEMBER √121 = 11 Also, (a+b)(a-b) = a^2 - b^2

OpenStudy (anonymous):

WOULD IT BE (11X+1)(11X-1)

OpenStudy (anonymous):

It would not. Did you see what I said above?

OpenStudy (anonymous):

(11x+3)(11x+3) = 121 x^2+66 x+9 AND (11x+3)(11x-3) = 121 x^2-9

OpenStudy (anonymous):

It is actually (-1+11 x)^2

OpenStudy (anonymous):

(11X+1)(11X-1)

OpenStudy (anonymous):

GOOD TRY THOUGH 121x^2 - 22x + 1 = (-1+11 x)(-1+11 x) = (-1+11 x)^2

OpenStudy (anonymous):

BUT ISNT IT THE SAME THIN I WROTE

OpenStudy (anonymous):

It isn't... You had one negative and one positive! Both are negative

OpenStudy (goformit100):

@romanortiz65 as you msged me, yes your answer is correct.

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

Your welcome?

OpenStudy (anonymous):

@goformit100 @romanortiz65 You see that answer is (a - b)^2, right?